Supplemental Operations Report – May 4th

Supplemental Operations Report 4 May 2022

Name of person filing report: Shannon Rupert

Reason for Report: May the fourth be with you! Happy Star Wars Day! Okay, so this is really just a routine report…

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: nothing to report

Generator: Still being manually run at night

ScienceDome Dual Split: Off

Solar—Charge controllers off at night

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 75 %

Propane Reading, director tank— 82 %

Propane Reading, intern tank— 70 %

Propane Reading, generator—61 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 500 gallons

Sojourner rover used: yes

Hours: 168.8

Beginning charge: 100

Ending charge: 64

Currently charging: no (will plug in tonight)

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Nothing to report

HabCar used and why, where? No

CrewCar used and why, where? Yes, to town.

Luna used and why, where? No

Campus-wide inspection, if action taken, what and why? Nothing to report

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Still working on the fence. It’s very hard to get t-posts into hardpan. The fence is now about half the size it was beyond the trailers.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: I need to go into the robotic observatory and take the cameras off the telescopes. Hoping to do this tomorrow.

Summary of health and safety issues: Nothing to report

EVA Report – May 4th

Crew 265 EVA Report 04-05-2022

EVA # 9

Author: Laude

Purpose of EVA: Social media filming by the Crew Journalist.

Start time: 0900
End time: 1140

Narrative: We made our way on foot to the base of Phobos Peak taking photos and videos along the way. Many interesting rock and sediment formations were seen with all displaying the forces of erosion likely by weather and water. We readily hiked a short way up to its southern slope and stopped at a nice nearly flat overlook, but could not view MDRS.

We then descended to the base and walked partway around the base looking for another perch. Finding none, we rehiked up to the previous overlook where Treadwell continued her work. We stayed there for about 40 minutes until she got everything she could think of and several minutes longer as Laude was reluctant to leave such an awesome view and time was on our side.

Finally, we made good speed and arrived at the hab 20 minutes early.

Destination: Phobos Peak

Coordinates (use UTM NAD27 CONUS): UTM NAD27 CONUS 519324 4250177

Participants: Laude and Treadwell

Road(s) and routes per MDRS Map: Walking to Phobos Peak.

Mode of travel: Foot

EVA Report – May 4th

Crew 265 EVA Report 04-05-2022

EVA # 10

Author: Levesque

Purpose of EVA: Long-range communications testing

Start time: 1300
End time: 1600

Narrative: The EVA team drove to Zubrin’s Head and then proceeded on foot to a point 500 meters west for a comms check. They returned to the rovers and drove down the road to repeat the process on foot on a road near White Rock Reservoir and later in White Rock Canyon to the south. They next drove back to the junction of Cow Dung Road and a rough road leading to Barrainca Butte and walked on foot from there to conduct two comm checks along that road. After completing all checks, they returned to the rovers and drove back to the Hab, having to stop twice because of a rover’s overheating.

Destination: Zubrin’s Head, White Rock Canyon, Barrainca Butte areas

Coordinates: UTM NAD27 CONUS 518804 4247962, 520100 4248538, 515968 4254298, 520369, 4247380, 519554 4247285, 518956 4247098

EVA Participants: Iakymov, Blanco, Licea

Road(s) and Routes per MDRS map: Cow Dung Road, Zubrin’s Head, White Rock Reservoir and Canyon, and Barrainca Butte.

Mode of Travel: Perserverance and Opportunity and on foot.

Sol Summary – May 4th

Crew 265 Sol Summary Report 04-05-2022

Sol: 10

Summary Title: Yeoman’s Service

Author’s Name: Marc Levesque, Crew Commander

Mission Status: Nominal

Sol Activity Summary:

The radio repeater battery finally called it quits after a cold night and yeoman’s service for nine sols, far beyond expectations.

Prior to the morning’s EVA departing, a discussion was held with the Station Director and crew members on how best to replace the battery, in sim or not. After some discussion, it was decided to quickly complete the replacement out of sim to maintain repeater operation, which was subsequently accomplished.

Following that, two EVAs were completed. Social media filming was done during a climb of Phobos Peak, and another EVA conducted comms checks south of the Hab in the White Rock Canyon and Barrainca Butte areas.

Look Ahead Plan: One EVA is planned for tomorrow and continued activity on mission projects. Work will also begin on end-of-mission reports, as the crew will be going out of sim on Friday morning.

Anomalies in work: None

Crew Physical Status: Good

EVA: None

Reports to be filed:

Sol Summary
Operations Report
HSO Report
Crew Journalist
EVA Request
EVA Report
Pictures and Picture of the Day

Support Requested: None

Operations Report – May 4th

Crew 265 Operations Report Sol 10 04-05-2022

SOL: 10

Name of person filing report: Dave Laude/Sergii Iakymov

Non-nominal systems: Furnace. Room#2 doorknob. Percy left rear tail light is non-functional but the brake lights working.

Notes on non-nominal systems: Furnace is functional and we remain cautious limiting to day use only. Doorknob feels like it’s about to fail.

Spirit rover used: No
Hours: 179.5
Beginning charge: 100
Ending charge: 100
Currently charging: Yes

Opportunity rover used: Yes
Hours: 97.6
Beginning charge: 100
Ending charge: 98
Currently charging: No

Curiosity rover used: No
Hours: 190.6
Beginning charge: 100
Ending charge: 100
Currently charging: Yes

Perseverance rover used: Yes
Hours: 239.6
Beginning charge: 100
Ending charge: 78
Currently charging: No

General notes and comments:

Summary of Hab operations:

Water (static tank): Not full. ~200 gallons.
Water (loft tank): 33 gallons
Water Meter: 158315 units

Static to Loft Pump used – Yes
Static tank pipe heater (on or off): Off
Static tank heater (On or off): Off
Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: Batteries charging.

Radio battery capacity testing has ended. 5 have >= 80% of new capacity that I call "Good" and 6 have >=60% and <80% that I call "Fair". None were <60%. Test results are on a label sticking to each battery.

For Scott Davis: I tried setting the one-piece suit #9 waist belt one notch lower (Lowest) and found it helpful in reducing shoulder weight to an acceptable level. Also, a good way to prevent over-discharging of the batteries and thus extending life is to use an electronic switch with an under-voltage cut-off with hysteresis for the fans (cutoff around 11.7V or ~ 90% discharged).

Summary of GreenHab operations: NA


Heater: N/A

Supplemental light: N/A

Harvest: None

Summary of ScienceDome operations: Not used.

Dual split: Heat or AC, On or Off

Summary of RAM operations: None

Summary of any observatory issues: NA

Summary of health and safety issues: All crew members are feeling good.

Questions, concerns and requests to Mission Support: Per Sergii’s calculation and comparison to the known static tank volume the water meter is pretty accurate. The meter is only 2.5% off on its readings from our calculated water usage.

Also, Dave has devised a method of more accurate water gauging of the static tank and verified by Sergii. It involves lowering a tape measure end down to just above the surface of the water. One will see the reflection of the tapes and easily know when to stop.

The distance from the top of the black rim to the surface is measured in meters and a simple formula is applied. It has been verified by calculation of a full tank to be 555 gallons, not to the rim, but to where the tank’s upper portion suddenly curves inward. Also, by use of the now known accurate meter.

Vol(gallons)=642-600H where H is the measure reading in meters (I know I’m mixing units of measure. It can be converted to any units).

It should be noted that the bottom 5cm (a guess) is not included as the pump will need some depth to work.

Journalist Report – May 4th

Crew 265 Journalist Report, May 4, 2022

Sol: 10

Summary Title: The End is Nigh…

Author: Sarah Treadwell, Crew Journalist

Sol 10. Two EVAs today, with myself and Executive Officer Dave in the morning and Engineer Sergii and Mapping Texas Twins Benny and Isai in the afternoon. We all sense the end of the mission drawing close, with one final EVA set for tomorrow. Overall, our mission seems to have been a success, but I will certainly let our crew commander do a final briefing on that.

This morning Dave and I attempted to summit nearby Phobos Peak. Unfortunately, there was no easy way to ascend as the ground was far too loose. After we went as high as we could go, we took some time to just sit and really soak in the environment around us. Those who know me well know that I can cry at a drop of a dime; however, today felt quite different. There was more of a sense of reverence that I carried with me, rather than overwhelming emotion.

The other EVA crew did some comms checking in a different direction from the station than we have normally been wandering. The radios seemed to perform well. They returned for us to all enjoy an early dinner and we parted ways into our sequestered quiet spaces to do homework and reports.

Generally, I think we all are just tired and ready to return to our normal lives… and even more importantly, return to non-dehydrated food. Personally, I am extremely excited about this and have been fantasizing about what to eat when we are done.

Something I want to coyly touch on that I will talk about more once I get home is one of my biggest mental challenges I have had while here. I think everyone can understand how not being able to call or video chat with our loved ones can be a challenge. While here I missed my daughter losing her first tooth, my boyfriend’s birthday, and my son getting to experience very sweet firsts at his preschool. While not ideal, these are all things that any astronaut or explorer has to accept they may miss when taking on a mission.

Something I did not anticipate though was receiving really, really big news and dealing with that in this environment where I couldn’t use my usual coping techniques to deal with my myriad of emotions. Normally I go for walks, drive to loud music, call my closest loved ones, and go to a restaurant to check out for an evening in a situation like this. None of those were options here.

This is something I will need time to reflect on because this is an important part of what we do here and what I also came to do, which is to look at the psychological and sociological effects of our time here.

But I will process that later… because all I can think about right now is having a burger and Coke. And so to end this, a quote for my son, whom we nickname Moo:

“Fix your little problem and light this candle” – Alan Shepherd

Copyright © The Mars Society. All rights reserved. | Main Site